NEW ANALYSIS OF THE CORIOLIS-INTERACTING ν_{2} AND ν_{5} BANDS OF CH_{3}^{79}Br AND CH_{3}^{81}Br

\underline{F. KWABIA TCHANA}, J. ORPHAL, I. KLEINER, CNRS, Laboratoire de Photophysique Moléculaire, 91405 Orsay Cedex, France; N. LACOME, Laboratoire de Dynamique, Interactions et Réactivité/Spectrochimie Moléculaire, Université Pierre et Marie Curie, UMR CNRS 7075, Case Courrier 49, 4 Place Jussieu, 75252 Paris Cedex 05, France; and O. BOUBA, Faculté des Sciences, Université de Ngaoundéré, B.P. 454 Ngaoundéré, Cameroun.

Methyl bromide (CH_{3}Br) has been identified as one of the major sources of atmospheric bromine. Atmospheric methyl bromide originates from both natural (algae, phytoplankton) and anthropogenic sources (agricultural fumigant): the tropospheric mixing ratio of CH_{3}Br is 9-11 pptV in the Northern hemisphere and about 8 pptv in the Southern hemisphere, with an increase of about 0.15 pptV per year^{a}. However, until present, no attempts have been made to determine atmospheric concentrations of CH_{3}Br using infrared spectroscopy. Although the line positions in this region have been studied previously at medium spectral resolution^{b}, little is known about the line intensities. The purpose of the present work is to complete and extend the previous studies of the 1200-1600 cm^{−1} spectral range and to provide a prediction of line positions and intensities accurate enough to determine optimal spectral windows for future atmospheric detection of CH_{3}Br. The ν_{2} (A_{1}) and ν_{5} (E) fundamental bands of CH_{3}^{79}Br and CH_{3}^{81}Br have been recorded at LPPM with a high-resolution Fourier-transform infrared spectrometer (unapodized resolution of 0.004 cm^{−1}). For both isotopomers, we assigned 3037 lines for the parallel bands, 4530 for the perpendicular bands, and in addition 80 perturbation-allowed transitions, with J ≤ 68 and K ≤ 11. By taking into account the xy-Coriolis interaction between the two bands, it has been possible to generate an accurate prediction of the whole spectrum, with a standard deviation of better than 7 × 10^{−4} cm^{−1}. The ground state axial rotational constants A_{0} were redetermined from allowed and perturbation-allowed infrared transitions observed in the ν_{2} and ν_{5} bands around the local crossing. The A_{0} values obtained for both isotopomers are more accurate but fully compatible with those obtained previously.

aWorld Meteorological Organization (WMO),“Scientific Assessment of Ozone Depletion: 1998”, Report 44, WMO Global Ozone Research And Monitoring Project, Chapter 2, Geneva, Switzerland, 1999.

bR. Anttila, C. Betrencourt-Stirnemann, and J. Dupre, J. Mol. Spectrosc., 100, 54, (1983).